40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13952     Entry time: Wed Jun 13 01:02:40 2018     In reply to: 13948     Reply to this: 13953   13966
Author: gautam 
Type: Update 
Category: LSC 
Subject: Reliable and repeatable 1f DRMI locking 

[koji, gautam]

With Koji's help, I got repeatable and reliable DRMI locking going again tonight - this is with the AS path optics for the spectroscopy measurement in place, although the AUX laser remained shuttered tonight. Results + spectra tomorrow, but here's what I did:

  • Initial alignment procedure was as usual - use arms+ASS to align ITMs, and then PRMI carrier+ASS to align PRM and BS.
  • Found the appropriate gains and demod phases.
  • Measured loop TFs - PRCL is a big mystery. Used these to finalize loop gains.
  • Ran some sensing lines.
  • Whitened DRMI PDs for a calibrated "low-noise" spectrum (although the coils were not de-whitened).

As I have found before, it is significantly easier to get the locking going post 11pm - the wall Seis BLRMS don't look that much quieter at midnight compared to 10pm, but this might be a scaling issue. I'll do a quantitative assessment next time... Also, Foton takes between 25-45 secs to save an updated filter (timed twice today).

ELOG V3.1.3-